OpenStudy (anonymous):

Help? Chocolate candy bars are on sale for $1.90 each at the supermarket. Jerome decides to take advantage of the sale and stock up. He brought $50 with him to the store, but he spends $3 on milk, $5 on a salad, and $2 on bread. What is the maximum number of chocolate candy bars he can buy?

OpenStudy (anonymous):

frist you add the prices of the milk, salad and bread which you get: 3 + 5 +2 = $10. so if you take that from $50, you have $40 left since: 50- 10 = 40. then you divide 1.90 from 40 where you get: 40 / 1.90 = 21.05....... then you round down to get 21 so the maximum number of chocolate candy bars he can buy is 21 :) i hope this is right and it helped you :)

OpenStudy (anonymous):

thanks! <3

OpenStudy (anonymous):

no problem :)